# Passing the Buck to AI: How Individuals’ Decision-Making Patterns Affect Reliance on AI

### v1

@@ Abstract.
Psychological research has identified different patterns individuals have while making decisions, such as vigilance (making decisions after thorough information gathering), hypervigilance (rushed and anxious decision-making), and buckpassing (deferring decisions to others). We examine whether these decision-making patterns shape peoples’ likelihood of seeking out or relying on AI. In an online experiment with 810 participants tasked with distinguishing food facts from myths, we found that a higher buckpassing tendency was positively correlated with both seeking out and relying on AI suggestions, while being negatively correlated with the time spent reading AI explanations. In contrast, the higher a participant tended towards vigilance, the more carefully they scrutinized the AI’s information, as indicated by an increased time spent looking through the AI’s explanations. These findings suggest that a person’s decision-making pattern plays a significant role in their adoption and reliance on AI, which provides a new understanding of individual differences in AI-assisted decision-making.
Human-AI Interaction, Decision-Making Pattern, AI-Assisted Decision Making, Information-seeking
ccs: Human-centered computing Empirical studies in HCIccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
To address this gap, this paper explores how individuals’ decision-making patterns impact whether people rely more on their own knowledge or on knowledge offered by an AI—in our case ChatGPT. When decisions have to be made, who seeks out the suggestions offered by ChatGPT? Who is more susceptible to this information despite the risks? And how are these choices related to individuals’ decision-making patterns?
@@
Our work builds on a well-established decision-making framework proposed by psychologists Irving Janis and Leon Mann (Mann et al., 1997), which has been widely validated with individuals from various countries (Cotrena et al., 2017; De Heredia et al., 2004; Colakkadioglu and Deniz, 2015; Bailly and Ilharragorry-Devaux, 2011; Nota et al., 2000). Mann et al. (1997) have found that people cope with decision-related stress differently, from vigilantly searching for information on their own to deferring decision-making to others. These decision-making patterns can be closely mapped to current interactions with AI systems: the amount of information individuals search during decision-making may impact how much information they need when evaluating AI suggestions; whether individuals defer decision-making to others in human-human interactions could influence their reliance on AI during human-AI interactions.
@@
To explore whether individual decision-making patterns predict a person’s AI-assisted decision-making tendencies, we preregistered111 See our preregistration on the Open Science Framework: https://osf.io/z9527/?view_only=ea63631787a74b3c8edd7a8ddc2dbd6b and conducted an online experiment (n=810) in which we asked participants to decide whether statements about food are facts or myths, letting them choose whether they wanted to seek ChatGPT’s decisions and explanations. To increase the stakes, participants were told to put together a pamphlet that should only include factual nutrition statements, and to imagine that the pamphlet will be shared with an organization that focuses on improving children’s health.
@@
Our results show that people differ in their interaction with ChatGPT depending on their individual decision-making pattern. Concretely, the higher a participants’ tendency to be a vigilant decision-maker, the more time they spent looking through ChatGPT’s explanations, underlining their tendency to scrutinize information before deciding. Conversely, the higher a participants’ tendency for buckpassing, the more likely they were to choose to see ChatGPT’s decisions, the less time they spent looking at its explanations for the decision, and the higher their self-reported reliance on ChatGPT. These findings indicate that, in general, buckpassers are less likely to question information provided by an AI, making them more susceptible to misinformation in AI responses than decision-makers who score low on buckpassing or high on vigilance.
@@
Overall, our work contributes (1) new empirical insights into how people differ in their interactions with AI, implying that parts of the population may be more vulnerable to the risks of AI; (2) design implications for AI tools that go beyond one-size-fits-all, including suggestions for leveraging the benefits of AI technologies while mitigating their risks; (3) a new, publicly available dataset (available at [URL omitted]) with the demographics, decision-making patterns, and AI interactions of 810 participants, which can be used by researchers for replicating and extending our analyses.
@@ 2.1. Studies of AI-assisted Decision-Making and Terminology
AI-assisted decision-making, also known as human-AI decision-making, broadly refers to scenarios where AI provides suggestions to help humans make decisions (Chen et al., 2023; Lai et al., 2023). To investigate the effectiveness of providing AI suggestions, recent research has examined users’ trust and reliance behaviors such as overall AI reliance, appropriate reliance, over-reliance, and under-reliance. Overall AI reliance measures users’ reliance behaviors in AI-assisted decision-making despite the desirability of such reliance. Users’ trust and reliance on systems are often captured via objective measures such as the agreement of an individual’s final decision and an AI’s suggestion in decision-making tasks (Papenmeier et al., 2022; Lu et al., 2024; Cao and Huang, 2022; Cao et al., 2023; Yin et al., 2019; Morrison et al., 2023) or switching from one’s decision to the suggestion of an AI (Yin et al., 2019; Morrison et al., 2023; Cao et al., 2023). Subjective measures such as perceived reliance and trust are also often used in prior studies (Cao and Huang, 2022; Buçinca et al., 2021). Generally, appropriate reliance is desired because it means that users only adopt the AI’s suggestions if they are correct. In contrast, over-reliance (when users adopt the AI suggestions despite it being inaccurate (Buçinca et al., 2021)) and under-reliance (where users fail to adopt the correct AI suggestions (Vasconcelos et al., 2023)) are undesirable. Studies examining these user behaviors often employ experimental designs in which AI suggestions are integrated into the collaboration process, either automatically presenting the AI suggestions before or after the users make their decisions (Chen et al., 2023; Morrison et al., 2023; Levy et al., 2021; Yin et al., 2019; Cao et al., 2023; Lu et al., 2024). As such, the measurement of reliance is often limited to this kind of experimental design, focusing on agreement with AI’s decision or switching. However, to examine whether AI suggestions are effective, it is also critical to understand whether and how people consider them, which might not be universal among individuals. Thus, our study extends the prior behavioral metrics for reliance and takes into account whether and how people choose to see AI suggestions.
@@ 2.2. Factors that Influence Trust and Reliance on AI
Recent studies have found a variance in individuals’ trust in, and reliance on AI suggestions and several factors related to system design and individual characteristics. In terms of system-related factors, evidence from empirical studies identified the effects of design and performance of AI systems such as the accuracy of the AI and whether the AI provides a reason for its decision (Lai and Tan, 2019; Wang and Yin, 2023; Buçinca et al., 2021; Schemmer et al., 2023; Schoeffer et al., 2024). Regarding user characteristics, evidence from survey studies suggests that users’ attitudes towards AI systems vary across demographics, including age, education levels, and knowledge about AI (Gillespie et al., 2023; Ma and Huo, 2023; Araujo et al., 2020; Chong et al., 2022). For example, in a study on individuals’ acceptance of ChatGPT, older individuals were found to be less willing to accept ChatGPT compared to younger individuals (Ma and Huo, 2023). Araujo et al. (2020) found that education levels and knowledge about AI and algorithms are positively associated with users’ perceived usefulness and fairness of AI. In addition, personal traits have also been found to play a role in users’ adoption of suggestions from AI systems: Chong et al. (2022), for instance, found that self-confidence directs individuals’ decisions to accept or reject suggestions from the AI. Focusing on individuals’ Big-Five personality traits (including openness, conscientiousness, agreeableness, neuroticism, and extraversion), Cai et al. (2022) found that individuals who scored high in conscientiousness—a tendency related to self-control and responsibility—have higher trust in AI systems that offer both human-requested and system-initiated suggestions.  Faruk et al. (2023) found students from India and Thailand who scored high in openness (being open to new experiences) are more likely to use ChatGPT. Meanwhile, they found negative effects of agreeableness (being cooperative, friendly) and neuroticism (high tendency toward negative feelings) on students’ usage of ChatGPT. Our study extends this line of work by focusing on the effects of individual decision-making patterns in AI-assisted decision-making.
@@ 2.3. Decision-Making Theories and Individual Patterns
Social science scholars have developed different theories explaining the mechanisms behind individuals’ decision-making. Rational Choice Theory, rooted in economics, posits that individuals calculate the utilities of all options to make decisions. Challenging this proposition, Simon (1955) has argued that people often seek satisfactory rather than optimal choices due to their limited capacity to evaluate numerous options. Schwartz et al. (2002) proposed that people tend to either have a tendency of maximizing or satisficing: maximizers tend to assess options carefully and choose one that provides maximum benefit later on, while satisficers tend to settle for an acceptable decision. Mann et al. (1997)’s theory takes into account individuals’ emotions and stress-coping patterns in their decision-making process. Their four decision-making patterns were derived from a survey study and are described as (i) vigilance (making decisions only after a comprehensive search of information), (ii) hypervigilance (approaching decisions in a hurried and anxious way), (iii) buckpassing (avoiding decisions or deferring to others), and (iv) procrastination (putting off the decision). See Table 1 for an overview and example scale items.
@@
According to Janis and Mann (1977), all four decision-making patterns are in the repertoire of individuals, but people tend to rely on one of these patterns more than on others. Importantly, Mann et al. (1997)’s decision-making patterns have been widely validated with participants from various countries and demographics (Cotrena et al., 2017; De Heredia et al., 2004; Colakkadioglu and Deniz, 2015; Bailly and Ilharragorry-Devaux, 2011; Nota et al., 2000) and have been shown to be relevant in real-world decision-making scenarios (Brown et al., 2016; Alexander et al., 2017; Kim et al., 2022). Moreover, Mann et al. (1998) revealed that individuals’ reliance on decision-making patterns (especially buckpassing) varies between East Asian and Western cultures. Together these studies suggest that people vary in their approach to decision-making.
@@
Decision-making patterns have been found to affect individuals’ decision outcomes. In the context of economic choices, research shows that when given a time constraint, maximizers tend to browse more options and change their decisions before making the final purchase than satisficers (Chowdhury et al., 2009). Adopting Mann et al. (1997)’s decision-making categories, Kim et al. (2022) found that individuals’ decision-making patterns are linked to their quality of life, including physical and psychological. These findings shed light on the impact of decision-making patterns on individuals’ lives, which could also be manifested in the context of AI-assisted decision-making. Jugovac et al. (2018) has examined how maximizers and satisficers differ in their interactions while providing AI recommendations, however, they found no effect. Our study extends this line of work by adopting the framework from Mann et al. (1998), studying how users’ behaviors differ when they are faced with uncertain information and have the option to defer to the AI.
@@ 3.1. Hypotheses
Overall, we hypothesize that decision-making patterns could affect individuals’ interactions with AI suggestions. Specifically, by AI suggestions, we refer to the recommended decisions from the AI (AI decisions) and its text-based justifications for its recommendation (AI explanations). Our study imagines scenarios where people encounter information online that may be right or wrong and use an AI tool like ChatGPT to assess its credibility. When using an AI, people can seek the AI’s decision and, if needed, the AI’s explanations for the decision. We speculate these AI suggestions are used differently across individuals with varying decision-making patterns, as outlined in Table 2. For vigilant decision-makers who tend to conduct a comprehensive search of information and evaluate it rigorously, we hypothesize that they are more likely to seek the AI’s suggestions for additional information even if they already have an answer. However, as they also tend to be more confident in their own decisions (Mann et al., 1998), they might not necessarily prioritize the AI’s decision over their own.
People who exhibit hypervigilant decision-making behaviors tend to search for information frantically, but not rigorously, before arriving at a decision. Thus, we hypothesize they will quickly seek information from the AI (i.e. see AI’s decision and explanations) without spending much time evaluating the information. As they experience high emotional stress and are less confident in their own decisions (Mann et al., 1998), they would be more likely to rely on the AI’s decision to reach their decisions.
@@
The higher people score on vigilance, the more likely they are to choose to see AI decisions. H1a
@@
The higher people score on hypervigilance, the more likely they are to choose to see AI decisions. H2a
@@
The higher people score on buckpassing, the more likely they are to choose to see AI decisions. H3a

### v2

@@ Abstract.
Psychological research has identified different patterns individuals have while making decisions, such as vigilance (making decisions after thorough information gathering), hypervigilance (rushed and anxious decision-making), and buckpassing (deferring decisions to others). We examine whether these decision-making patterns affect peoples’ engagement with AI-generated information in decision-making. In an online experiment with 810 participants tasked with distinguishing food facts from myths, we found that a higher buckpassing tendency was positively correlated with the likelihood of seeking AI information and reported reliance on AI, while being negatively correlated with the time spent reading AI explanations. In contrast, the higher a participant tended towards vigilance, the more carefully they scrutinized the AI’s information, as indicated by an increased time spent looking through the AI’s explanations. These findings suggest that a person’s decision-making pattern plays a significant role in their interactions with AI suggestions, which provides a new understanding of individual differences in AI-assisted decision-making.
decision-making, user characteristics, psychology, Human-AI interaction
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
To address this gap, this paper explores how individuals’ decision-making patterns impact their tendency to rely on seeing knowledge offered by an AI—in our case ChatGPT. As such, we conceptualize individuals’ reliance on AI as their engagement with AI-generated information when AI suggestions are available. When decisions have to be made, who seeks out the suggestions offered by ChatGPT? Who is more susceptible to this information despite the risks? And how are these choices related to individuals’ decision-making patterns?
@@
Our work builds on a well-established decision-making framework proposed by psychologists Irving Janis and Leon Mann (Mann et al., 1997), which has been widely validated with individuals from various countries (Cotrena et al., 2017; De Heredia et al., 2004; Colakkadioglu and Deniz, 2015; Bailly and Ilharragorry-Devaux, 2011; Nota et al., 2000). Under this framework, decision-making patterns are coping patterns that correspond to situational stress. Adoption of these patterns varies among individuals since people cope with decision-related stress differently, from vigilantly searching for information on their own to deferring decision-making to others. These decision-making patterns can be closely mapped to current interactions with AI systems: the amount of information individuals search during decision-making may impact how much information they need when evaluating AI suggestions; whether individuals defer decision-making to others in human-human interactions could influence their reliance on AI during human-AI interactions.
@@
To explore whether individual decision-making patterns predict a person’s AI-assisted decision-making tendencies, we preregistered [^1] See our preregistration on the Open Science Framework: https://osf.io/z9527/?view_only=ea63631787a74b3c8edd7a8ddc2dbd6b and conducted an online experiment (n=810) in which we asked participants to decide whether statements about food are facts or myths, letting them choose whether they wanted to seek ChatGPT’s decisions and explanations. To increase the stakes, participants were told to put together a pamphlet that should only include factual nutrition statements, and to imagine that the pamphlet will be shared with an organization that focuses on improving children’s health.
@@
Our results show that people differ in their interaction with ChatGPT depending on their individual decision-making pattern. Concretely, the higher a participant’s tendency to be a vigilant decision-maker, the more time they spent looking through ChatGPT’s explanations, underlining their tendency to scrutinize information before deciding. Conversely, the higher a participant’s tendency for buckpassing, the more likely they were to choose to see ChatGPT’s decisions, the less time they spent looking at its explanations for the decision, and the higher their self-reported reliance on ChatGPT. These findings indicate that, in general, buckpassers are less likely to question information provided by an AI, making them more susceptible to misinformation in AI responses than decision-makers who score low on buckpassing or high on vigilance.
@@
Overall, our work contributes (1) new empirical insights into how people differ in their interactions with AI, implying that parts of the population may be more vulnerable to the risks of AI; (2) design implications for AI tools that go beyond one-size-fits-all, including suggestions for leveraging the benefits of AI technologies while mitigating their risks; (3) a new, publicly available dataset (available at the GitHub repository [^2] https://github.com/Mooniem/passing_the_buck_to_AI) with the demographics, decision-making patterns, and AI interactions of 810 participants, which can be used by researchers for replicating and extending our analyses.
@@ 2.1. Studies of AI-assisted Decision-Making and Terminology
AI-assisted decision-making, also known as human-AI decision-making, broadly refers to scenarios where AI provides suggestions to help humans make decisions (Chen et al., 2023; Lai et al., 2023). To investigate the effectiveness of providing AI suggestions, recent research has examined users’ trust and reliance behaviors such as overall AI reliance, appropriate reliance, over-reliance, and under-reliance. Overall AI reliance measures users’ reliance behaviors in AI-assisted decision-making regardless of the desirability of such reliance. Users’ trust and reliance on systems are often captured via objective measures such as the agreement of an individual’s final decision and an AI’s suggestion in decision-making tasks (Papenmeier et al., 2022; Lu et al., 2024; Cao and Huang, 2022; Cao et al., 2023; Yin et al., 2019; Morrison et al., 2023) or switching from one’s decision to the suggestion of an AI (Yin et al., 2019; Morrison et al., 2023; Cao et al., 2023). While behavioral metrics were also used to quantify individuals’ trust in AI systems (Yin et al., 2019), later researchers pointed out that trust and reliance are distinct concepts with different properties (Lee and See, 2004). Trust is considered to guide yet not necessarily determine reliance behaviors. Thus these behavioral metrics are often used to measure reliance, and separate trust-related questionnaires are used to measure trust in AI systems. Subjective measures such as perceived reliance are also often used in prior studies (Cao and Huang, 2022; Buçinca et al., 2021). Generally, appropriate reliance is desired because it means that users only adopt the AI’s suggestions if they are correct. In contrast, over-reliance (when users adopt the AI suggestions despite it being inaccurate (Buçinca et al., 2021)) and under-reliance (where users fail to adopt the correct AI suggestions (Vasconcelos et al., 2023)) are undesirable. Notably, studies examining these user behaviors often employ a two-stage decision-making workflow in which AI suggestions are integrated into the collaboration process, by automatically presenting the AI suggestions before or after the users make their decisions (Chen et al., 2023; Morrison et al., 2023; Levy et al., 2021; Yin et al., 2019; Cao et al., 2023; Lu et al., 2024). As such, the measurement of reliance is often limited to the experimental design that focuses on agreement with AI’s decision or switching. However, to examine whether AI suggestions are effective, it is also critical to understand whether and how people consider them, which might not be universal among individuals. Thus, our study extends the prior behavioral metrics for reliance and takes into account whether and how people choose to see AI suggestions.
@@ 2.2. Factors that Influence Trust and Reliance on AI
Recent studies find variance in individuals’ trust in, and reliance on AI suggestions and several factors related to system design and individual characteristics. In terms of system-related factors, evidence from empirical studies identified the effects of design and performance of AI systems such as the accuracy of the AI and whether the AI provides a reason for its decision (Lai and Tan, 2019; Wang and Yin, 2023; Buçinca et al., 2021; Schemmer et al., 2023; Schoeffer et al., 2024). Regarding user characteristics, evidence from survey studies suggests that users’ attitudes towards AI systems vary across demographics, including age, education levels, and knowledge about AI (Gillespie et al., 2023; Ma and Huo, 2023; Araujo et al., 2020; Chong et al., 2022). For example, in a study on individuals’ acceptance of ChatGPT, older individuals are found to be less willing to accept ChatGPT compared to younger individuals (Ma and Huo, 2023). Araujo et al. (2020) find that education levels and knowledge about AI and algorithms are positively associated with users’ perceived usefulness and fairness of AI. In addition, personal traits have also been found to play a role in users’ adoption of suggestions from AI systems: Chong et al. (2022), for instance, find that self-confidence directs individuals’ decisions to accept or reject suggestions from the AI. Focusing on individuals’ Big-Five personality traits (including openness, conscientiousness, agreeableness, neuroticism, and extraversion), Cai et al. (2022) identify that individuals who score high in conscientiousness—a tendency related to self-control and responsibility—have higher trust in AI systems that offer both human-requested and system-initiated suggestions. Faruk et al. (2023) show that students from India and Thailand who score high in openness (being open to new experiences) are more likely to use ChatGPT. Meanwhile, they find negative effects of agreeableness (being cooperative, friendly) and neuroticism (high tendency toward negative feelings) on students’ usage of ChatGPT. Our study extends this line of work by focusing on the effects of individual decision-making patterns in AI-assisted decision-making. While decision-making patterns are found to relate to decision self-esteem and have points of contact with individual differences on personality dimensions such as self-efficacy (Mann et al., 1997), they hold distinct value in human decision-making theory. Unlike personality dimensions that are stable traits, these patterns are individuals’ coping strategies with situational stress, which can be influenced by context, offering value for design interventions that help adapt to individuals’ decision states.
@@ 2.3. Decision-Making Theories and Individual Patterns
Scholars from various social science disciplines—such as economics and psychology—have developed different theories explaining the mechanisms behind individuals’ decision-making. The rational choice theory, rooted in economics, posits that individuals calculate the utilities of all options to make decisions. Challenging this proposition, Simon (1955) argues that people often seek satisfactory rather than optimal choices due to their limited capacity to evaluate numerous options. Schwartz et al. (2002) propose that people tend to either have a tendency of maximizing or satisficing: maximizers tend to assess options carefully and choose one that provides maximum benefit later on, while satisficers tend to settle for an acceptable decision. Different from the economists’ perspectives that focus on the utility of decisions, psychologists develop constructs and models of decision-making that relate to motivational processes, cognitive styles, and personality dimensions and traits. The motivational processes refer to decision-making processes in response to certain motivations such as preserving self-image or dealing with negative thoughts and feelings. Example constructs for cognitive styles include need for cognition (the tendency to exert greater cognitive effort to evaluate information) (Cacioppo and Petty, 1982). Mann et al. (1997)’s theory takes into account individuals’ emotions and stress-coping patterns (motivational processes) in their decision-making process. Their four decision-making patterns are derived from a survey study and are described as (i) vigilance (making decisions only after a comprehensive search of information), (ii) hypervigilance (approaching decisions in a hurried and anxious way), (iii) buckpassing (avoiding decisions or deferring to others), and (iv) procrastination (putting off the decision). See Table 1 for an overview and example scale items.
@@
According to Janis and Mann (1977), all four decision-making patterns are in the repertoire of individuals, but people tend to rely on one of these patterns more than on others. Importantly, Mann et al. (1997)’s decision-making patterns have been widely validated with participants from various countries and demographics (Cotrena et al., 2017; De Heredia et al., 2004; Colakkadioglu and Deniz, 2015; Bailly and Ilharragorry-Devaux, 2011; Nota et al., 2000) and have been shown to be relevant in real-world decision-making scenarios (Brown et al., 2016; Alexander et al., 2017; Kim et al., 2022). Moreover, Mann et al. (1998) reveal that individuals’ reliance on decision-making patterns (especially buckpassing) varies between East Asian and Western cultures. Together these studies suggest that people vary in their approach to decision-making. They also found that individuals’ decision self-esteem is positively correlated with vigilance and negatively with hypervigilance and buckpassing.
@@
Decision-making patterns have been found to affect individuals’ decision outcomes. In the context of economic choices, research shows that when given a time constraint, maximizers tend to browse more options and change their decisions before making the final purchase than satisficers (Chowdhury et al., 2009). Adopting Mann et al. (1997)’s decision-making categories, Kim et al. (2022) found that individuals’ decision-making patterns are linked to their quality of life, including physical and psychological. These findings shed light on the impact of decision-making patterns on individuals’ lives, which could also be manifested in the context of AI-assisted decision-making. Jugovac et al. (2018) has examined how maximizers and satisficers differ in their interactions while presenting AI recommendations; however, they found no effect. Notably, their study design focuses on participants’ interactions with two provided recommendations instead of a design setting where relying on an AI is optional. Our study further explores this line of work in a design setting where AI support is optional, examining how users’ behaviors differ when they are faced with uncertain information and have the option to defer to the AI.
@@ 3.1. Hypotheses
[…]
Overall, we hypothesize that decision-making patterns could affect individuals’ interactions with AI suggestions. Specifically, by AI suggestions, we refer to the recommended decisions from the AI (AI decisions) and its text-based justifications for its recommendation (AI explanations). Our study imagines scenarios where people encounter information online that may be right or wrong and use an AI tool (e.g., ChatGPT) to assess its credibility. When using an AI, people can seek the AI’s decision and, if needed, the AI’s explanations for the decision. We speculate these AI suggestions are used differently across individuals with varying decision-making patterns, as outlined in Table 2. For vigilant decision-makers who tend to conduct a comprehensive search of information and evaluate it rigorously, we hypothesize that they are more likely to seek the AI’s suggestions for additional information even if they already have an answer. However, as they also tend to be more confident in their own decisions (Mann et al., 1998), they might not necessarily prioritize the AI’s decision over their own.
People who exhibit hypervigilant decision-making behaviors tend to search for information frantically, but not rigorously, before arriving at a decision. Thus, we hypothesize they will quickly seek information from the AI (i.e., see AI decision and explanations) without spending much time evaluating the information. As they experience high emotional stress and are less confident in their decisions (Mann et al., 1998), they would be more likely to rely on the AI’s suggestions to reach their decisions.
@@
The higher people score on vigilance, the more likely they are to choose to see AI decision. H1a
@@
The higher people score on hypervigilance, the more likely they are to choose to see AI decision. H2a
@@
The higher people score on buckpassing, the more likely they are to choose to see AI decision. H3a